A cricket ball is thrown vertically up with a speed of 20 m/s from ground. Find the speed of the ball when it has reached a height of 15 m.


1
Expert's answer
2021-05-27T08:31:25-0400

According to the conservation energy law, the current energy of the ball (which is the sum of its kinetic and potential energies) is equal to its initial kinetic energy:


"\dfrac{mv_0^2}{2} = \dfrac{mv^2}{2} + mgh"

where "m" is the mass of the ball, "v_0 = 20m\/s" is its initial speed, "h =15m" is its current height, "v" is its current speed, "g = 9.8 m\/s^2" is the gravitational acceleration. Expressing "v", obtain:


"v = \sqrt{v_0^2 - 2gh}\\
v = \sqrt{20^2 - 2\cdot 9.8\cdot 15} \approx 10m/s"

Answer. 10 m/s.